OpenStudy (anonymous):

A student scored 71,83 and 96 on three tests. What must he score on the fourth test in order to have an average of 85?

OpenStudy (anonymous):

(71 + 83 + 96 + x) / 4 = 85 340 = 71 + 83 + 96 + x solve for x x = 90